A boy is now two-thirds as old as his sister, and two years ago he was five-eighths as old as she was then. Find their present ages.
Answer by stanbon(75887)

A boy is now two-thirds as old as his sister, and two years ago he was five-eighths as old as she was then. Find their present ages.
---------------
Equations:
b = (2/3)s
b-2 = (5/8)(s-2)
---------------------
Substitute and solve for "s"::
(2/3)s - 2 = (5/8)s - (10/8)
----------------
Multiply thru by 24 to get:
16s - 48 = 15s - 30
----
s = 78 yrs (sister's age now)
----
Solve for "b":
b = (2/3)s
b = (2/3)78
b = 52 yrs (boy's age now)
------------
Cheers,
Stan H.